x-3y+2z = 6
2x-y+3z=5
z + y -z = -4

Answer by ikleyn(52794)

.
x-3y+2z = 6
2x-y+3z=5
z + y -z = -4
~~~~~~~~~~~~~~~~~~~~~

From your third equation y = -4 (if the equation is written correctly).